OpenStudy (anonymous):

The hypotenuse of an isosceles right triangle is 7 feet longer than either of its legs. Find the exact length of each side.

OpenStudy (anonymous):

r u still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{a^2+a^2}=\sqrt{2a^2}=\sqrt{2}a\] a is length of one of the leg

OpenStudy (anonymous):

how do i do it?

OpenStudy (anonymous):

hypotenuse=\[\sqrt{2}a\] hypotenuse=7+a

OpenStudy (anonymous):

7+a=sqrt{2}a

OpenStudy (anonymous):

Can you draw out the whole equation?

OpenStudy (anonymous):

Draw the triangle. Label the hypotenuse as a+7, and the side as a. Then use the pythagorean theorem (since its a right triangle), to relate the two as abdul did.

OpenStudy (anonymous):

Can you draw it?